A and B started a business with 1500 and 2500 and got a profit 800 rs. Half of the profit is shared equally the remaining is shared according to their investment. Find their profits.

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Hemant

Remaining is 400

A will take = (1500/1500+2500)*400 = (1500/4000)*400 = 150

B will take 400-150=250

A=200+150=350

B=200+250=450

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..